75cc of water sample requires 10.2 cc of 0.01 M EDTA for titration in presence of
eriochrome black T indicator. Calculate the total hardness.
1
Expert's answer
2020-02-27T09:08:29-0500

M1V1=M2V2

M1=M2V2/V1= 10.2 x 0.01 / 75=0.00136 M.

0.00136 x 10^5 = 136 ppm